# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Coding

### v1

@@ 3.2. Participant identification
We began by searching GitHub to identify repositories related to code generation via AI. We did this by developing a list of relevant tags, including “coding-assistant” and “ai-code-generator.” The full list of tags can be found in our replication package (ano, 2024). This resulted in 27 repositories, including open-source coding assistants such as “Tabby” (tabby, [n. d.]) and “cptX” (cptX, [n. d.]) as well as “full development” solutions such as “gpt-engineer” (gptengineer, [n. d.]) and “gpt-pilot” (gpt-pilot, [n. d.]). We included an additional three repositories we believed would be relevant: Meta’s CodeLlama, (B. Rozière et al., 2024), GitHub Copilot (GitHub, [n. d.]), and Codeium (codeium, [n. d.]). This brought our total repository count to 30. Next, we sought to identify developers who had contributed to or shown interest in any of these repositories. To do this, we used the GitHub API (GitHub REST API documentation, [n. d.]) to collect account information for users that forked, starred, watched, or contributed to any of the 30 aforementioned repositories. Lastly, in order to respect the privacy of developers, we filtered this information to eliminate users who did not provide publicly available contact information on their GitHub profile. This resulted in 30,659 public-facing email addresses, which we used to contact individuals requesting participation in our research study. All information reported in this paper, including demographic information, was contributed voluntarily by respondents who agreed to participate in response to our request.
@@ 3.3. Survey response collection and analysis
Survey responses to five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Two SE researchers performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 574 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. Our replication package (ano, 2024) contains the final codes and definitions.
@@
We did not base our analysis on inter-rater agreements because multiple codes could be assigned to each response, and no list of codes existed before the start of coding (i.e., our approach was fundamentally inductive). However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the results’ reliability. All of the responses were coded in a single iteration by the two annotators. To mitigate agreement by chance, the labels assigned to each answer were reviewed by the annotators, including those without disagreement. We avoided defining “umbrella” codes by creating complete definitions for each code and sharing them between the annotators during review, as well as by assessing the appropriateness of codes during reconciliation of the annotators’ results such that codes that were too similar could be merged together and codes that were overbroad could be broken into multiple, more specific codes. This was done when the annotators observed multiple codes frequently being used to describe the same statement in a given answer or when a code was being applied to answers with significantly different meanings, respectively.
@@
Two open-ended questions solicited additional thoughts, comments, or explanations on respondents’ answers to a number of different questions. The disparity in responses prevented meaningful analysis with our previous coding method, so we applied a different strategy. One SE researcher read through all 574 responses for both questions and grouped them into categories and sub-categories based on their content by considering the respondent’s other answers as context. An additional researcher reviewed and validated these categorization decisions, discussing and resolving disagreements when needed. Lastly, a third author reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For the closed-ended questions, we aggregated results using descriptive statistics and discussed them to resolve any disagreements. One author manually coded each response for questions with an “Other” option (such as those asking about languages or tools used). Sections 4, 5, and 6 present the results of our analysis.
@@ 3.4. Participant demographics and background
Respondents self-selected the nature of their development work from a list that included proprietary software development (331), open source (247), and academia (117). The total number of responses to this question is greater than the number of complete and valid responses (574) because respondents could select more than one option. Respondents came from six continents, as indicated in Figure 4, and 73 different countries, the top ten of which are shown in Table 3.
Of the 574 complete and valid responses we obtained, 554 developers (96.5%) indicated that they were generally familiar with using GenAI for producing source code. Of those, 497 (89.7%) had incorporated such tools into their development workflow, as indicated in Table 1. Respondents also used a wide array of programming languages, 48 in total. The most popular languages used were Python, JavaScript, C/C++, and Java, as indicated in Table 2. Respondents had experience with many different AI tools, including GitHub Copilot (73.6%), ChatGPT (84.9%), Bard/Gemini (19.1%), and Claude (10.3%), as well as several open-source models (7.8%); more than 60 tools and models were reported in total, as indicated in Table 4.
@@ 3.5. Follow-up Interviews
All survey participants were asked at the conclusion of the survey if they would be willing to be contacted for a follow-up interview. Those who consented by supplying their email address (382, 66.5%) made up our potential interview participant pool. Two authors then selected 22 potential interview candidates from this pool based on various criteria, including experience, geographic jurisdiction, and indicated views/perceptions on key legal issues, with the goal of interviewing participants representing a diversity of views. The full list of interview selection criteria can be found in our replication package (ano, 2024). Of the 22 candidates, seven responded to schedule an interview. The seven interviewees represented four continents (North America (3), Europe (2), South America (1), and Africa (1)) and a variety of backgrounds in software development, including working as a developer on proprietary software (6), contributing to OSS (4), and working in academia (1).
@@
Each interview was conducted jointly by two SE researchers, and the recording and transcript for each interview was independently reviewed by one interviewer to ensure the accuracy of the transcript and extract the relevant portions of each response, assigning topic labels to each one. These quotes and topic labels were hosted in a shared document accessible to both researchers, allowing the reuse of topic labels. Each researcher analyzed roughly half of the interviews. Afterwards, the two researchers jointly reviewed the labels assigned to each response to verify the accuracy and completeness of the analysis, discussing and reaching a consensus on the framing and application of topic labels. A third researcher reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For simplicity and clarity, when we refer to quotes from follow-up interviews, we use the notation RI and describe participants by their assigned survey IDs.
@@
4. RQ1 Use of Generative AI technologies for software development
RQ1 aimed to discover how developers are using GenAI technologies for developing software. Although the developers surveyed had a range of perceptions about AI and copyright issues, most seemed to share the belief that the use of GenAI is now a regular part of everyday development activities. As respondent $R_{90}$ commented, “[D]evs not using AI is like accountants not using Excel in the 80s.” In a follow-up interview, $RI_{149}$ commented, “If one of my engineers isn’t using a code generation model, I’m going to fire him pretty soon. Because there’s very little virtue in doing useless grunt work. […] I actually expect it to be used on every part of the project. If it isn’t, that’s the exception. In fact, I want to know why [it wasn’t used].” In what follows, we present and discuss the findings for this RQ. A summary of key findings can be found in Table 5.
@@ 4.1. Uses, benefits, and challenges of GenAI
Before we can properly understand developers’ thoughts and perspectives on copyright and other legal issues pertaining to GenAI, we must first explore the context in which they use these tools. In this section, we consider the usage scenarios identified by developers, the benefits they have derived from this use, and the challenges that they face.
@@ 4.1.1. GenAI usage in software development
Almost all respondents (554) stated that they were generally familiar with the use of GenAI to produce source code. Of these, the vast majority—497 (89.7%)—indicated that they had in some capacity incorporated such tools into their development workflows. The remaining 57 (10.3%) offered several reasons for not regularly using GenAI, including legal risk (8), personal preference (8), the lack of significant use cases warranting its use (6), reluctance to send sensitive information to a remote server (6), constraints imposed by an employer (6), and limited trust in the tooling (5).
Among those who had a personal preference against using GenAI, some developers conveyed the sense that GenAI encourages monotony or lack of effort. This was somewhat surprising because, ideally, GenAI is used to automate repetitive and boring tasks, as previous work has indicated (Sergeyuk et al., 2024). $R_{188}$ stated “I like to think for myself. [W]ithout [that], coding becomes boring and a lot of reading instead of reading and writing.” $R_{339}$ put it this way: “It’s better to write code on my own than to review code generated by AI.” $R_{276}$ reported similarly, “I’ve tested incorporating GitHub’s [C]opilot into my workflow and caught myself many times waiting for the autocompletion to kick in so I could just press ‘Tab’ and go from there. Soon I realized I was getting lazy and not thinking much, which I didn’t enjoy.” The full list of reasons put forward by developers can be found in our replication package (ano, 2024).
Finding 1: While the vast majority of developers surveyed (89.7%) were using GenAI tools, there were many reasons why developers might not incorporate GenAI into their development workflows, including perceived legal issues, personal preference, information security concerns, and limitations placed by employers.
Of those familiar with GenAI tools, relatively few developers (32) said that their organization did not permit the use of GenAI for coding. These developers reported their understanding of the restriction. Twelve (37.5%) stated that such usage could compromise sensitive or proprietary information. Ten (31.3%) indicated that their organizations were concerned about the legal implications of using such tools. Four (12.5%) suggested that their employers were slow to adopt new technology and did not fully understand its benefits. Three respondents (9.4%) pointed to the fear of security issues, such as incorporating insecure code or designs. Notably, these responses could reflect overlapping concerns: compromising proprietary information, for example, could also raise legal concerns.
Finding 2: Of those respondents who reported working at organizations that disallowed the use of GenAI, many indicated that they believed the restriction was based on fear of proprietary information leakage and the potential legal challenges such usage might bring. Some developers also perceived that their employers were slow to adopt new technology.
@@
Respondents have used GenAI for a variety of software development tasks (as seen in Figure 5), including writing source code (427), writing code documentation (309), test case generation (279), refactoring (268), prototyping (251), creating UIs (152), and code styling/formatting (146). Developers also mentioned that they used such tools for debugging (12), reference or learning (12), code explanation (8), troubleshooting (6), repository management (5), and brainstorming (4).
@@ 4.1.2. GenAI tools used by developers
A wide variety of GenAI tools, including open-source, open-weight, and proprietary models, were used by participants. In total, developers reported using 64 distinct, named GenAI tools, the top 20 of which can be seen in Table 4. Developers were encouraged to identify all the tools that they had used in their work. Of these, the top six (which were also provided in a list of choices, along with an “Other” option) were: GitHub Copilot (366), ChatGPT-4 (304), ChatGPT-3.5 (268), Bard (95), Claude (51), and Amazon CodeWhisperer (33). The next largest group (19) reported that they used open-source/open-weight models, with several such tools/models listed by multiple respondents, including Codellama (11), Mistral (9), Ollama (7), Llama2 (7), and Mixtral (6). Developers also used other proprietary models and services, including Codeium (16), Perplexity (12), Phind (10), and JetBrains AI (7), and six respondents indicated that they used custom-trained or in-house models for their development work. A full list of the tools developers reported using can be found in our replication package (ano, 2024).
Of the tools listed by respondents, 20 were designed to be integrated into an IDE. Many of these tools provide autocomplete suggestions for developers based on the context of surrounding code, but some also include Q/A chat functionality. Developers also reported using 18 distinct web-based tools. These tools, such as ChatGPT (OpenAI, [n. d.]), Perplexity (Perplexity, [n. d.]), and Hugging Face Chat (HuggingChat, [n. d.]), offer great convienence, but require the user to supply more context. Respondents also indicated using four different model infrastructures: Ollama (Ollama, [n. d.]), Oobabooga (Oobabooga, [n. d.]), llama.cpp (Llama.cpp, [n. d.]), and Mozilla llamafile (Hood, 2023). These infrastructures provide a streamlined means to run and train open-weight models locally. In total, 15 distinct open-weight models/model families were also identified by respondents. (We note that the prevalence of open-weight models and open-source tools in this space has likely increased since our initial survey with the subsequent release of models such as Llama3 (Meta, 2024) and the open-sourcing of XAI’s Grok (XAI, 2024).) Some developers also noted the use of GenAI agents like gpt-engineer (gptengineer, [n. d.]) and metagpt (Hong et al., 2023), which allow models to self-prompt with the goal of accomplishing more complex tasks. Lastly, four GenAI tools which were not strictly for software development, such as Figma (figma, [n. d.]) and GrammarlyGo (grammarly, [n. d.]), were mentioned.
@@
Figure 6 shows the reported tool usages across the categories previously discussed. Since developers selected all the tools they had used for development tasks and many of those tools might fall into the same categories, we can’t generalize to reach a conclusion about any developer preferences. We can also say little about the frequencies with which each tool type is used. That said, we see that there were 761 reported usages of web-based tools. There were only 464 reported usages of IDE-based tools, despite them constituting the largest number of distinct tools. This might suggest that developers were more likely to experiment with various different web-based tooling solutions, especially during the early days of GenAI for software development. There were 74 reported usages of open-weight or locally hosted models, but only 11 instances of reported corresponding model infrastructure. It could be that many developers using open-weight models neglected to mention the tooling they used to run them or that they relied on web-hosted solutions such as Hugging Face Chat.
Finding 3: The ecosystem of GenAI tools used by developers was varied and extensive, comprising both open-source/open-weight and proprietary solutions. Developers primarily made use of web-based tools and those that could be integrated into their IDEs.
@@ 4.1.3. Benefits derived from GenAI usage
Respondents who had used GenAI as part of their workflow identified several benefits from its use. The greatest benefit mentioned by 288 developers (58%) was increased productivity, faster development, and efficiency. Other benefits included not having to spend time writing simple or boilerplate code (88), quick prototyping (42), debugging (37), faster, more accurate documentation (29), summarizing existing documentation as an alternative to reading it (28), providing inspiration (16), refactoring (14), use in writing test cases (5), domain logic support (4), and generating test data (4). A full list of the benefits identified by respondents can be found in our replication package (ano, 2024).
Thirty respondents identified improvements in code quality and optimization over what human developers could achieve. $RI_{118}$ noted, “When you couple the lack of training [of the average developer] with generally one to three years’ experience, and the fact is that GPT4 or many of these tools are actually better developers than they are. […] I would almost feel better if [a code comment] said I wrote this and then I got a code review from GPT4.” $R_{74}$ stated that GenAI’s ability to “notic[e] optimizations or cleaner ways of writing something like a custom function is quite useful.” Some respondents, like $R_{466}$, noted that the “[generated] code [is] more reusable and better,” and $R_{669}$ noted that the output of code generation models was “following best [coding] practices.”
Similarly, developers related that GenAI could help them write software in unfamiliar languages (28) and libraries (27). During a follow-up interview, $RI_{190}$ described how they would “basically just writ[e] the code in Typescript and then just tell [ChatGPT] convert to Liquid” and that “otherwise [they] would probably quit because Liquid is incredibly painful.” Using GenAI specifically in the capacity of porting or translating code was mentioned by six respondents. $R_{286}$ found that GenAI “was helpful to migrate code from a middleware platform to Javascript based development.” $R_{90}$ summarized this benefit: “The syntax of various languages is [why it] can take hundreds of hours of use to establish [familiarity] in a new language, but AI significantly reduces the ramp-up time.”
Along the same lines, several respondents noted that GenAI helped them to be better developers by enhancing their understanding of unfamiliar code. For example, nineteen respondents stated that GenAI provided “assistance in [the] understanding of code snippets” ($R_{47}$). $R_{404}$ also explained that “[GenAI] is also very good at providing semi-formal abstract explanations of the topics at hand and it provides consistent analogies which makes things very smooth to follow and comprehend.” But GenAI models are not limited to explaining source code or concepts. $R_{238}$ noted that they could also be used for “commit message generation [and ex]planation of changes in commit (summarization).”
Finding 4: The benefits developers reported from using GenAI tools for coding go beyond increasing productivity and avoiding repetitive, boring tasks. They also included understanding code better and improved code quality and code optimization, which some human developers might struggle to achieve.

### v3

@@ 3.2. Participant identification
We began by searching GitHub to identify repositories related to code generation via AI. We did this by developing a list of relevant tags, including “coding-assistant” and “ai-code-generator.” The full list of tags can be found in our replication package (ano, 2024). This resulted in 27 repositories, including open-source coding assistants such as “Tabby” (tabby, [n. d.]) and “cptX” (cptX, [n. d.]) as well as “full development” solutions such as “gpt-engineer” (gptengineer, [n. d.]) and “gpt-pilot” (gpt-pilot, [n. d.]). We included an additional three repositories we believed would be relevant: Meta’s CodeLlama, (B. Rozière et al., 2024), GitHub Copilot (GitHub, [n. d.]), and Codeium (codeium, [n. d.]). This brought our total repository count to 30. Next, using the GitHub API (GitHub REST API documentation, [n. d.]), we sought to identify developers who had shown interest in any of these repositories by forking, starring, watching, or contributing to them. We believe that these actions are highly correlated with familiarity with using GenAI tools for software development tasks, although we also acknowledge that our respondent sample may not be representative of all developers and may overrepresent those who frequently use GenAI tools (see Section 7). Lastly, to respect the privacy of developers, we filtered the collected information to eliminate users who did not provide publicly available contact information on their GitHub profile. This resulted in 30,659 public-facing email addresses, which we used to contact individuals seeking participation in our research study. All information reported in this paper, including demographic information, was contributed voluntarily by respondents who agreed to participate in response to our request.
@@ 3.3. Survey response collection and analysis
Survey responses to five open-ended questions were analyzed through a qualitative coding approach (Spencer, 2009). Two authors, both SE researchers, performed open-coding by independently assigning one or more codes to each response using a shared spreadsheet and codebook. Each annotator independently coded all 574 responses, adding new codes to the codebook as necessary. Once the initial coding was completed, the annotators met to settle disagreements and consolidate the set of codes. The disagreements encountered consisted of differing interpretations of the response, applying the same codes in different contexts, or using broader versus more specific codes. Through this process, all such disagreements were resolved, resulting in a final set of codes, definitions, and assignments for each response. Our replication package (ano, 2024) contains the final codes and definitions.
@@
We did not base our analysis on inter-rater agreements because multiple codes could be assigned to each response, and no list of codes existed before the start of coding (i.e., our approach was fundamentally inductive). However, we carefully followed best open-coding practices (Spencer, 2009), and we leveraged coders’ discussions to ensure the results’ reliability. All of the responses were coded in a single iteration by the two annotators. To mitigate agreement by chance and the introduction of bias, the labels assigned to each answer were reviewed by the annotators, including those without disagreement. The same annotators performed this review so as to be able to provide their understanding and rationale behind the labeling of each response. We avoided defining “umbrella” codes by creating complete definitions for each code and sharing them between the annotators during the review, as well as by assessing the appropriateness of codes during reconciliation of the annotators’ results such that codes that were too similar could be merged together and codes that were overbroad could be broken into multiple, more specific codes. This was done when the annotators observed multiple codes frequently being used to describe the same statement in a given answer or when a code was being applied to answers with significantly different meanings, respectively.
@@
Two open-ended questions solicited additional thoughts, comments, or explanations on respondents’ answers to a number of different questions. The disparity in responses prevented meaningful analysis with our previous coding method, so we applied a different strategy. One SE researcher read through all 574 responses for both questions and grouped them into categories and sub-categories based on their content by considering the respondent’s other answers as context. An additional researcher reviewed and validated these categorization decisions, discussing and resolving disagreements when needed. Lastly, as a separate task, a third author reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For the closed-ended questions, we aggregated results using descriptive statistics and discussed them to resolve any disagreements. One author manually coded each response for questions that had an “Other” option (such as C2 and D4), so that this additional information could be effectively analyzed. Sections 4, 5, and 6 present the results of our analysis.
The law researcher contributed an expert understanding of the legal context relating to software licensing issues. This perspective facilitated the interpretation of the survey results and the discussion of the applicable landscape and legal risks presented by the technology.
@@ 3.4. Participant demographics and background
Continent Count Perc.
North America 198 34.49%
Europe 163 28.4%
Asia 159 27.7%
Africa 27 4.7%
South America 21 3.66%
Oceania 15 2.61%
Unknown 13 2.26%
Table 4. Continent of Origin*
@@
Respondents self-selected the nature of their development work from a list that included proprietary software development (331), open source (247), and academia (117). The total number of responses to this question is greater than the number of complete and valid responses (574) because respondents could select more than one option. Respondents came from six continents, as indicated in Table 4, and 73 different countries, the top ten of which are shown in Table 3.
Of the 574 complete and valid responses we obtained, 554 developers (96.5%) indicated that they were generally familiar with using GenAI for producing source code. Of those, 497 (89.7%) had incorporated GenAI tools into their development workflow, as indicated in Table 1. Respondents also used a wide array of programming languages, 48 in total. The most popular languages used were Python, JavaScript, C/C++, and Java, as indicated in Table 2. Respondents had experience with many different AI tools, including GitHub Copilot (73.6%), ChatGPT (84.9%), Bard/Gemini (19.1%), and Claude (10.3%), as well as several open-source models (7.8%); more than 60 tools and models were reported in total, the top 20 of which are indicated in Table 5.
@@ 3.5. Follow-up Interviews
All survey participants were asked at the conclusion of the survey if they would be willing to be contacted for a follow-up interview. Those who consented by supplying their email address (382, 66.5%) made up our potential interview participant pool. Two authors then selected 22 potential interview candidates from this pool based on various criteria, including experience, likely legal jurisdiction based on location, and indicated views/perceptions on key legal issues, with the goal of interviewing participants representing a diversity of views. The full list of interview selection criteria can be found in our replication package (ano, 2024). Of the 22 candidates, seven responded to schedule an interview. The seven interviewees represented four continents (North America (3), Europe (2), South America (1), and Africa (1)) and a variety of backgrounds in software development, including working as a developer on proprietary software (6), contributing to OSS (4), and working in academia (1).
@@
Each interview was conducted jointly by two authors, both SE researchers. The same researchers conducted all of the interviews. The recording and transcript for each interview were independently reviewed by one of the two interviewers to ensure the accuracy of the transcript and extract the relevant portions of each response, assigning topic labels to each one. These quotes and topic labels were hosted in a shared document accessible to both researchers, allowing the reuse of topic labels. Each researcher analyzed roughly half of the interviews. Afterward, the two researchers jointly reviewed the labels assigned to each response to verify the accuracy and completeness of the analysis, discussing and reaching a consensus on the framing and application of topic labels. A third researcher reviewed the quotes extracted for the paper, making sure none were miscategorized, misattributed, or taken out of context. For simplicity and clarity, when we refer to quotes from follow-up interviews, we use the notation RI and describe participants by their assigned survey IDs.
@@
4. RQ1: Use of Generative AI technologies for software development
RQ1 aimed to discover how developers are using GenAI technologies for developing software. Although the developers surveyed had a range of perceptions about AI and copyright issues, most seemed to share the belief that the use of GenAI is now a regular part of everyday development activities.
In the following we present and discuss the findings for this RQ. We present the key conclusions that answer this research question as Findings 1-10. A summary of key findings can be found in our replication package (ano, 2024).
@@ 4.1. Uses, benefits, and challenges of GenAI
In this section, we consider the usage scenarios identified by developers, the benefits developers have derived from this use, and the challenges that developers face.
@@ 4.1.1. GenAI usage in software development
Almost all respondents (554) stated that they were generally familiar with the use of GenAI to produce source code. Of these, the vast majority—497 (90%)—indicated that they had in some capacity incorporated such tools into their development workflows. (Figure 3 displays the types of tasks mentioned by respondents.) The remaining 57 (10%) offered several reasons for not regularly using GenAI, including legal risk (8), personal preference (8), the lack of significant use cases warranting its use (6), reluctance to send sensitive information to a remote server (6), constraints imposed by an employer (6), and limited trust in the tooling (5).
Among those who had a personal preference against using GenAI, some developers conveyed through open-ended responses the sense that GenAI encourages monotony or lack of effort, despite the fact that GenAI is often used to automate repetitive and boring tasks, as previous work has indicated (Sergeyuk et al., 2024). $R_{188}$ stated, “I like to think for myself. [W]ithout [that], coding becomes boring and a lot of reading instead of reading and writing.” $R_{339}$ put it this way: “It’s better to write code on my own than to review code generated by AI.” The full list of reasons put forward by developers can be found in our replication package (ano, 2024).
Finding 1: While the vast majority of developers we surveyed (90%) were using GenAI tools, there were many reasons why some developers did not incorporate GenAI into their development workflows, including perceived legal issues, personal preference, information security concerns, and limitations placed by employers.
Of those familiar with GenAI tools, relatively few developers (32) said that their organization did not permit the use of GenAI for coding. These developers reported their understanding of the restriction. Twelve (38%) stated that such usage could compromise sensitive or proprietary information. Ten (31%) indicated that their organizations were concerned about the legal implications of using such tools. Four (13%) suggested that their employers were slow to adopt new technology and did not fully understand its benefits. Three respondents (9%) pointed to the fear of security issues, such as incorporating insecure code or designs. Notably, these responses could reflect overlapping concerns: compromising proprietary information, for example, could also raise legal concerns.
Finding 2: Of the relatively few respondents who reported working at organizations that disallowed the use of GenAI, about a third indicated that they believed the restriction was based on fear of proprietary information leakage and the potential legal challenges that such usage might bring. Some developers also perceived that their employers were slow to adopt new technology.
@@ 4.1.2. GenAI tools used by developers
Participants used various GenAI tools, including open-source, open-weight, and proprietary models. In total, developers reported using 64 distinct, named GenAI tools, the top 20 of which can be seen in Table 5. Developers were encouraged to identify all the tools they had used in their work. Of these, the top six (which were also provided in a list of choices, along with an “Other” option) were: GitHub Copilot (366), ChatGPT-4 (304), ChatGPT-3.5 (268), Bard (95), Claude (51), and Amazon CodeWhisperer (33). The next largest group (19) reported that they used open-source/open-weight models, with several such tools/models listed by multiple respondents, including Codellama (11), Mistral (9), Ollama (7), Llama2 (7), and Mixtral (6). Developers also used other proprietary models and services, including Codeium (16), Perplexity (12), Phind (10), and JetBrains AI (7), and six respondents indicated that they used custom-trained or in-house models for their development work. A full list of the tools developers reported using can be found in our replication package (ano, 2024).
Of the tools listed by respondents, 20 were designed to be integrated into an IDE. Many of these tools provide autocomplete suggestions for developers based on the context of surrounding code, but some also include Q/A chat functionality. Developers also reported using 18 distinct web-based tools, such as ChatGPT, Perplexity, and Hugging Face Chat (OpenAI, [n. d.]; Perplexity, [n. d.]; HuggingChat, [n. d.]), that require the user to supply more context than IDE-integrated tools. Respondents also indicated using four model infrastructures: Ollama, Oobabooga, llama.cpp, and Mozilla llamafile (Ollama, [n. d.]; Oobabooga, [n. d.]; Llama.cpp, [n. d.]; Hood, 2023). These infrastructures provide a streamlined means to run and train open-weight models locally. In total, respondents identified 15 distinct open-weight models/model families. (We note that the prevalence of open-weight models and open-source tools in this space has likely increased since our initial survey with the subsequent release of models such as Llama3 (Meta, 2024) and the open-sourcing of XAI’s Grok (XAI, 2024).) Three developers also noted the use of GenAI agents like gpt-engineer (gptengineer, [n. d.]) and metagpt (Hong et al., 2023), which allow models to self-prompt to accomplish more complex tasks. Lastly, four GenAI tools that were not designed for software development tasks, such as Figma (figma, [n. d.]) and GrammarlyGo (grammarly, [n. d.]), were mentioned. Figure 4 shows the usage of tools in the categories discussed here as reported by respondents.
@@
Finding 3: The ecosystem of GenAI tools used by developers was varied and extensive, comprising both open-source/open-weight and proprietary solutions. Developers primarily used web-based tools and those that could be integrated into their IDEs.
@@ 4.1.3. Benefits derived from GenAI usage
Respondents who had used GenAI as part of their workflow identified several benefits from its use. The greatest benefit mentioned by 288 developers (58%) was increased productivity, faster development, and efficiency. Other benefits included not having to spend time writing simple or boilerplate code (88), quick prototyping (42), debugging (37), faster, more accurate documentation (29), summarizing existing documentation as an alternative to reading it (28), providing inspiration (16), refactoring (14), use in writing test cases (5), domain logic support (4), and generating test data (4). A full list of the benefits identified by respondents can be found in our replication package (ano, 2024). We note that participants often mentioned tasks as benefits, likely indicating that GenAI improved the effectiveness, productivity, or ease of completion for those particular tasks.
Thirty respondents identified improvements in code quality and optimization over what human developers could achieve. $RI_{118}$ noted, “When you couple the lack of training [of the average developer] with generally one to three years’ experience, […] the fact is that GPT4 or many of these tools are actually better developers than they are. […] I would almost feel better if [a code comment] said I wrote this and then I got a code review from GPT4.” Similarly, developers related that GenAI could help them write software in unfamiliar languages (28) and libraries (27). During a follow-up interview, $RI_{190}$ described how they would “basically just writ[e] the code in Typescript and then just tell [ChatGPT] convert to Liquid.” Six respondents mentioned using GenAI specifically to port or translate code. $R_{90}$ summarized this benefit: “The syntax of various languages is [why it] can take hundreds of hours of use to establish [familiarity] in a new language, but AI significantly reduces the ramp-up time.”
Along the same lines, nineteen respondents noted that GenAI helped them to be better developers by enhancing their understanding of unfamiliar code. However, GenAI models are not limited to explaining source code or concepts. $R_{238}$ noted that they could also be used for “commit message generation [and ex]planation of changes in commit (summarization).”
Finding 4: The benefits developers reported from using GenAI tools for coding go beyond increasing productivity and avoiding repetitive, boring tasks. They also included a better understanding of code, improved code quality, and code optimization, which some human developers might struggle to achieve.